What is the market value of a 20-year bond with face value of $1000, which makes quarterly coupon payments at a coupon rate of 10%, if the required rate of return is 8% per year? Round off your final answer to three digits after the decimal point.
Answer #1

value of a bond is present value of future cash flows discounted at required return

coupon rate = 10%

so rate per period = 10 / 4 = 2.5% (quartelry return)

coupon per period = 1000 x 2.5% = 25

required return per period = 8 / 4 = 2%

maturity period = 20 years

number of payments = 20 x 4 = 80 (80 periods)

market value:

fir --PV(B4,B2,B7,B1) A 1 Maturity(Par) value 1,000 2 Number of periods untill maturity 3 Coupon rate per period 2.50% 4 yiel

so market value of the bond = $1,198.723

(formula is given in above sheet)